OpenStudy (anonymous):

The bank is willing loan you money to purchase a home at 1% interest, compounded monthly. If you would like to buy a $300000 house, how much would your payments be each month be if you take out a: 10 year mortgage? 20 year mortgage? 30 year mortgage? Do not round your answers!

OpenStudy (lilg132):

300,000 / 100 = 3000 ( This is 1% of 300,000 ) ( This is also the interest every month) now if you want a 10 year mortgage you need to find out how much you will be paying back so you take the value of the house and divide it by 10 years to get the amount you will have to pay each year then divide it by 12 to get how much you will have to pay every month plus the interest you will have to pay every month 300,000 / 10 = 30,000 (This is how much you will have to pay per year) 30,000 / 12 = 2500 2500 per month are just repayments for the house but you need to also include the 1% interest which is 3000 so 2500 + 3000 = $5500 per month on a 10 year mortgage now we leave the amount of interest the same on each of the mortgages but the price we have to pay back will be different as we have more time so to work out how much we will have to pay back on a 20 year mortgage, take the value of the house divide it by 20 and then divide it by 12 300,000 / 20 = 15,000 15,000 / 12 = 1250 now you add this to the 1% interest which is 3000 1250 + 3000 = $4250 per month on a 20 year mortgage and you do the same thing for the 30 year mortgage 300,000 / 30 = 10,000 10,000 / 12 = 833.3333333333333 833.3333333333333 + 3000 = $3833.3333333333333 every month on a 30 year mortgage dont forget to click good answer :)